OpenStudy (anonymous):

A height above ground, h metres, of a child on a swing, ranges from 0.5m to 1.5m The child reaches max height ever second. Initially there child is at minimum height. Express height as a function of time. I get h(t)=0.5sin(1t)+1, is this correct?

ganeshie8 (ganeshie8):

Very close! try this : \[\large h(t)=0.5\sin(2\pi t)+1\]

ganeshie8 (ganeshie8):

that doesn't look correct either.. looks we need to slide the graph to right because the starting value is minimum

ganeshie8 (ganeshie8):

` Initially there child is at minimum height.`

ganeshie8 (ganeshie8):

try this instead : \[\large y=0.5\sin\left(2\pi t-\frac{\pi}{2}\right)+1\]

OpenStudy (anonymous):

why 2pi t?

ganeshie8 (ganeshie8):

because you want the period to be `1` : `The child reaches max height ever second.`

ganeshie8 (ganeshie8):

the period of sin(t) is 2pi since you want to reduce the period to 1, multiplying t by 2pi compresses the graph by a factor of 2pi

OpenStudy (anonymous):

ah ok :)

ganeshie8 (ganeshie8):

see the graph https://www.desmos.com/calculator/gxe9rljmmy

OpenStudy (anonymous):

Thank you so very very much :)

ganeshie8 (ganeshie8):

yw!